Problem 1 [10 pts] A 20.0 kg child plays on a swing having support ropes that are 2.0 m long. A friend pulls her back until the ropes are 45.0° from the vertical and releases her from rest.

a) What is the potential energy for the child just as she is released, compared with the potential energy at the bottom of the swing?

b) How fast will she be moving at the bottom of the swing?

c) How much work does the tension in the ropes do as the child swings from the initial position to the bottom?
Answer #1

(a) Potential energy N= mgh - 0 Here m-mass g = grauitastion constant = 9.8 m/s2 h = the vertical difference of when Height w2x 114.856 m 2x114.858 v² = 229.712 ( v2 = 11.485 v= J11-485 = 3.388 m/ Work (W) = F.d F = mg = 20 kg x 9.8 m/s2 = 196N Here